# UDPNet: Unleashing Depth-based Priors for Robust Image Dehazing

### v1

@@ 5.3 Comparisons with SOTA Methods
We first present a comprehensive comparison with recent state-of-the-art approaches on the SOTS dataset Li et al (2018) in Table 2. As shown, our proposed methods consistently achieve superior performance across both datasets. In particular, ConvIR + UDP obtains a remarkable PSNR of 43.12 dB on SOTS-Indoor, surpassing the previous best ConvIR-B Cui et al (2024a) and MB-TaylorFormerV2-L Jin et al (2025) by 0.40 dB and 0.28 dB, respectively. On the SOTS-Outdoor dataset, it also achieves a competitive 40.32 dB PSNR, outperforming most recent models such as C2PNet Zheng et al (2023), and DEA-Net Chen et al (2024). More importantly, our FSNet + UDP further pushes the performance boundary, setting a new state-of-the-art with 43.30 dB PSNR on SOTS-Indoor, while maintaining comparable results on SOTS-Outdoor. These results highlight the effectiveness of our UDP design, achieving a new milestone in image dehazing. Moreover, we further evaluate our approach on the more realistic synthetic Haze4K dataset Liu et al (2021a). Table 3 showcases the performance of our UDP method applied to FSNet. As reported in Table 3, applying our UDP strategy to FSNet yields an average PSNR improvement of 1.19 dB over the original model. In addition, ConvIR + UDP also achieves a competitive PSNR of 34.82 dB. The qualitative comparisons on these daytime datasets, including SOTS-Indoor Li et al (2018), SOTS-Outdoor Li et al (2018), and Haze4K Liu et al (2021a), are illustrated in Figs. 7, 7, and 8, respectively. Our models generate haze-free images that are more visually consistent with the ground truth.
@@
To further demonstrate the superiority of our model, we conduct experiments on two nighttime datasets, i.e., GTA5 Yan et al (2020) and NHR Zhang et al (2020). Table 4 shows that ConvIR + UDP and PoolNet + UDP secure the first and second places with 33.12 dB and 32.78 dB PSNR, surpassing the original counterparts by 1.29 dB and 1.25 dB, respectively. The original models struggle to achieve optimal results under low-light conditions. However, by incorporating our UDP approach, we effectively alleviate these limitations, enhancing the overall brightness and recovering fine image details. The visual comparisons in Fig. 9 further confirm the effectiveness of our design, as the images generated by our models are much closer to the ground truth. We further extensively compare our models with state-of-the-art methods on real-world datasets. As shown in Table 6, our methods achieve state-of-the-art performance on both Dense-Haze Ancuti et al (2019) and NH-HAZE Ancuti et al (2020). On Dense-Haze, FSNet + UDP obtains 17.85 dB PSNR, surpassing FSNet by 0.72 dB. On NH-HAZE, ConvIR + UDP achieves the best performance with 20.98 dB PSNR, outperforming the previous best MB-TaylorFormerV2-B Jin et al (2025), and also attains the best perceptual quality. These results clearly demonstrate the effectiveness of our UDP in boosting real-world dehazing performance. The visual comparisons are shown in Fig. 11.
@@
Since image dehazing is crucial for remote sensing applications, we further evaluate our models on the SateHaze1k Huang et al (2020) dataset in Table 7. Our methods consistently achieve the best results across all three subsets. In particular, PoolNet + UDP attains 26.20 dB, 28.26 dB, and 23.13 dB PSNR on the three subsets, outperforming both general dehazing models (e.g., FocalNet Cui et al (2023a), ConvIR-S Cui et al (2024b)) and remote sensing–oriented designs (e.g., Trinity-Net Chi et al (2023)). Notably, on the challenging Thick subset, our method achieves a gain of 1.44 dB over FocalNet, highlighting its strong generalization capability to remote sensing scenarios.
@@
To further demonstrate the robustness and superiority of our approach, we conduct experiments on the all-in-one setting Jiang et al (2025). This fundamental setting evaluates the ability of a single model to function as a universal and versatile restorer. In our study, we adopt the more challenging five-task benchmark, which jointly considers Gaussian denoising, deraining, dehazing, motion deblurring, and low-light enhancement, thus requiring stronger generalization from a single set of model weights. As shown in Table 8, both PromptIR + UDP and AdaIR + UDP achieve substantial improvements in dehazing and low-light enhancement, with PromptIR + UDP yields further gains in deraining and AdaIR + UDP further boosting performance in motion deblurring. This can be largely attributed to the integration of depth-based priors, which promote a more nuanced understanding of global atmospheric illumination and fine-grained object structures. Although originally introduced to enhance dehazing, such priors also strengthen the model’s ability to recover structural fidelity in motion deblurring and illumination details in low-light enhancement. Consequently, AdaIR + UDP surpasses its original counterpart and establishes a new benchmark for all-in-one image restoration.
@@
Table 8: Performance comparison of different methods on five challenging benchmarks. Denoising results are reported at $\sigma=25$. The baseline results are in blue. Best results are highlighted.
@@
  Method Dehazing on SOTS Deraining on Rain100L Denoising on BSD68 Deblurring on GoPro Low-Light on LOL Average
Li et al (2018) Yang et al (2019) Martin et al (2001) Nah et al (2017) Wei et al (2018)
@@
  (TIP’23) DehazeFormer Song et al (2023) 25.31 0.937 33.68 0.954 30.89 0.880 25.93 0.785 21.31 0.819 27.42 0.875
(ICCV’23) Retinexformer Cai et al (2023) 24.81 0.933 32.68 0.940 30.84 0.880 25.09 0.779 22.76 0.863 27.24 0.873
(ICCVW’21) SwinIR Liang et al (2021) 21.50 0.891 30.78 0.923 30.59 0.868 24.52 0.773 17.81 0.723 25.04 0.835
(CVPR’22) DGUNet Mou et al (2022) 24.78 0.940 36.62 0.971 31.10 0.883 27.25 0.837 21.87 0.823 28.32 0.891
(CVPR’22) Restormer Zamir et al (2022) 24.09 0.927 34.81 0.960 31.49 0.884 27.22 0.829 20.41 0.806 27.60 0.881
(ECCV’22) NAFNet Chen et al (2022) 25.23 0.939 35.56 0.967 31.02 0.883 26.53 0.808 20.49 0.809 27.76 0.881
(TPAMI’23) FSNet Cui et al (2023b) 25.53 0.943 36.07 0.968 31.33 0.883 28.32 0.869 22.29 0.829 28.71 0.898
(TPAMI’19) DL Fan et al (2019) 20.54 0.826 21.96 0.762 23.09 0.745 19.86 0.672 19.83 0.712 21.05 0.743
(CVPR’22) Transweather Valanarasu et al (2022) 21.32 0.885 29.43 0.905 29.00 0.841 25.12 0.757 21.21 0.792 25.22 0.836
(CVPR’22) AirNet Li et al (2022) 21.04 0.884 32.98 0.951 30.91 0.882 24.35 0.781 18.18 0.735 25.49 0.846
(NeurIPS’23) PromptIR Potlapalli et al (2023) 26.54 0.949 36.37 0.970 31.47 0.886 28.71 0.881 22.68 0.832 29.15 0.904
(ICLR’25) AdaIR Cui et al (2025b) 30.53 0.978 38.02 0.981 31.35 0.889 28.12 0.858 23.00 0.845 30.20 0.910
(ICLR’25) $\text{DCPT}_{\text{PromptIR}}$ Hu et al (2025) 30.72 0.977 37.32 0.978 31.32 0.885 28.84 0.877 23.35 0.840 30.31 0.911
(TPAMI’25) DA-RCOT Tang et al (2025) 30.96 0.975 37.87 0.980 31.23 0.888 28.68 0.872 23.25 0.836 30.40 0.911
(TIP’25) Perceive-IR Zhang et al (2025) 28.19 0.964 37.25 0.977 31.44 0.887 29.46 0.886 22.88 0.833 29.84 0.909
(TIP’25) Pool-AIO Cui et al (2025a) 30.25 0.977 37.85 0.981 31.24 0.887 27.66 0.844 22.66 0.841 29.93 0.906
(TIP’25) $\text{DPPD}_{\text{PromptIR}}$ Wu et al (2025) 30.31 0.980 37.32 0.980 31.33 0.885 28.74 0.875 22.73 0.846 30.09 0.913
(CVPR’25) VLU-Net Zeng et al (2025) 30.84 0.980 38.54 0.982 31.43 0.891 27.46 0.840 22.29 0.833 30.11 0.905
@@
Table 9: Ablation analysis on different variants of our method, mainly including two aspects: fusion stages and attention types. Here, all ablation models adopt the backbone of FSNet Cui et al (2023b).
@@
5.4 Ablation Study
We conduct ablation studies to demonstrate the effectiveness of the proposed UDPNet by training and testing the models on the Haze4K dataset Liu et al (2021a). The model is trained for 1000 epochs using FSNet Cui et al (2023b) as the baseline.
5.4.1 Contribution of the Depth Map
The first step in our analysis is to justify the integration of depth information. As shown in Table 10, the baseline model FSNet achieves a PSNR of 34.12 dB. By incorporating depth maps extracted using the DepthAnything-V2-small variant together with our proposed UDP, the model’s performance is improved to 34.69 dB, while only introducing 0.32 M parameters. Although the gain at this stage is relatively modest, it demonstrates the effectiveness of depth priors. Using the base and large variants of DepthAnything V2 further increases performance, reaching 34.87 dB and a peak of 35.31 dB, respectively. In contrast, when the depth map is replaced with a 128-level gray map, the performance drops to 34.08 dB, highlighting that the performance improvement comes from accurate depth information rather than the increase in parameter count. Overall, these results confirm that accurate depth estimation is a key factor in image dehazing.
Table 10: Comparison of different depth maps. All depth variants have an identical parameter increase of only +0.32M.
@@ 5.4.1 Contribution of the Depth Map
Figure 13: Performance and efficiency variations with our UDP on the Haze4K dataset, compared with baseline methods Cui et al (2023b, 2024a, 2025a).
5.4.2 Effectiveness of each module
We perform the breakdown ablations by applying our modules to the baseline successively. The number of heads in the multi-head attention is set to 1, while all other configurations remain identical to the final dehazing model. As shown in Table 9, directly concatenating refined depth maps and RGB images at the network head (row (b)) significantly improves performance compared to the baseline (row (a)), reaching 34.78 dB PSNR. This demonstrates the substantial benefit of introducing depth maps for image dehazing. By contrast, replacing simple concatenation with the proposed DGAM (row (c)) yields further gains with 34.96 dB PSNR, verifying that adaptively weighting depth cues through channel attention is more effective than naive feature stacking.
We further explore two variants of DPFM: channel cross-attention (CCA) and spatial cross-attention (SCA). Results in rows (d–g) show that both variants outperform DGAM, indicating the benefit of explicitly modeling cross-modal dependencies. Notably, using RGB features as the query (CCArgb, SCArgb) consistently outperforms the depth-as-query counterparts, suggesting that RGB semantics provide a more reliable guidance signal. Moreover, spatial cross-attention (SCA) shows slightly higher gains than channel cross-attention (CCA), suggesting that spatial correspondences play a crucial role in depth–RGB fusion. Across all cases, the best performance 35.22 dB PSNR is achieved in row (l), where both SCAdepth and SCArgb are combined in the encoder, highlighting the complementarity of depth and RGB at the spatial level and confirming the necessity of deliberate design for depth map integration.
@@ 5.4.2 Effectiveness of each module
5.4.3 Position of DPFM
We also investigate inserting DPFM into different network stages. Placing DPFM in the encoder (rows (d–g)) yields more notable gains in Table 9, since incorporating depth priors at early stages enhances semantic abstraction and provides stronger guidance for subsequent layers. By contrast, applying DPFM to the decoder (rows (h–k)) still brings consistent improvements, but the benefits are relatively smaller. This can be attributed to the fact that depth maps mainly contain low-frequency structural cues, which may interfere with the reconstruction of high-frequency image details when introduced in the decoder.
5.4.4 Impact of the number of heads in DPFM
@@ 5.4.5 Generalization of our proposed UDP
To evaluate the adaptability of the proposed UDP, we integrate it into different baseline models, including FSNet Cui et al (2023b), ConvIR Cui et al (2024a), PoolNet Cui et al (2025a), PromptIR Potlapalli et al (2023) and AdaIR Cui et al (2025b). Each modified network is then trained and tested under the same experimental settings. The comparative results demonstrate that UDP can not only be applied to a variety of multi-scale image restoration networks but also effectively enhance the performance of all-in-one image restoration models, yielding improvements in tasks such as dehazing, deblurring, deraining, and low-light enhancement.
5.5 Limitations
@@
Our work introduces a novel method for image dehazing by effectively leveraging depth-based priors, which significantly enhance model adaptability and performance through plug-and-play modules. By integrating depth features, our approach lays the foundation for developing more robust and versatile dehazing models. However, the UDP framework has several limitations that open avenues for future research. One key limitation is the reliance on depth priors extracted from degraded images, which may introduce inaccuracies due to noise or occlusions in the depth estimation process. These inaccuracies could lead to structural distortions or loss of fine details in the restored image. Furthermore, the dependence on DepthAnything V2 priors during inference introduces computational inefficiency, as the external model requires additional resources for depth estimation, increasing inference time. Future work could explore alternative strategies for real-time depth inference or distilling depth priors from depth estimation models into image restoration models. While the UDP framework demonstrates promising results across various restoration tasks, the potential of leveraging depth-based priors for general image restoration remains an exciting direction. Future studies will focus on refining adaptive strategies to improve performance in a broader range of image restoration scenarios, with particular attention to enhancing real-time performance and reducing reliance on external depth models.
6 Conclusion

### v2

@@ V-C Comparisons with SOTA Methods
We first present a comprehensive comparison with recent approaches on the SOTS dataset [li2018benchmarking] in Table II. As shown, our proposed methods consistently achieve superior performance across both datasets. In particular, ConvIR + UDP obtains a remarkable PSNR of 43.12 dB on SOTS-Indoor, surpassing the previous best ConvIR-B [cui2024revitalizing] and MB-TaylorFormerV2-L [jin2025mb] by 0.40 dB and 0.28 dB, respectively. On the SOTS-Outdoor dataset, it also achieves a competitive 40.32 dB PSNR, outperforming most recent models such as C2PNet [zheng2023curricular], and DEA-Net [chen2024dea]. More importantly, our FSNet + UDP further pushes the performance boundary, setting a new state-of-the-art with 43.30 dB PSNR on SOTS-Indoor and 40.53 dB PSNR on SOTS-Outdoor. These results highlight the effectiveness of our UDP design, achieving a new milestone in image dehazing. Moreover, we further evaluate our approach on the more realistic synthetic Haze4K dataset [liu2021synthetic]. Table III showcases the performance of our UDP method applied to FSNet. As reported in Table III, applying our UDP to FSNet yields an average PSNR improvement of 1.19 dB over the original model. In addition, ConvIR + UDP also achieves a competitive PSNR of 34.82 dB. The qualitative comparisons on these daytime datasets, are illustrated in Figs. 6 and 7, respectively. Our models generate haze-free images that are more visually consistent with the GT.
@@
To further demonstrate the superiority of our model, we conduct experiments on two nighttime datasets, i.e., GTA5 [yan2020nighttime] and NHR [zhang2020nighttime]. Table IV shows that ConvIR + UDP and PoolNet + UDP secure the first and second places with 33.12 dB and 32.78 dB PSNR, surpassing the original counterparts by 1.29 dB and 1.25 dB, respectively. The original models struggle to achieve optimal results under low-light conditions. However, by incorporating our UDP approach, we effectively alleviate these limitations, enhancing the overall brightness and recovering fine image details. The visual comparisons in Fig. 8 further confirm the effectiveness of our design, as the images generated by our models are much closer to the ground truth. We further extensively compare our models with state-of-the-art methods on real-world datasets. As shown in Table V, our methods achieve state-of-the-art performance on both Dense-Haze [ancuti2019dense] and NH-HAZE [ancuti2020nh]. On Dense-Haze, FSNet + UDP obtains 17.85 dB PSNR, surpassing FSNet by 0.72 dB. On NH-HAZE, ConvIR + UDP achieves the best performance with 20.98 dB PSNR, outperforming the previous best MB-TaylorFormerV2-B [jin2025mb], and also attains the best perceptual quality. These results clearly demonstrate the effectiveness of our UDP in boosting real-world dehazing performance. The visual comparisons are shown in Fig. 10.
@@
Image dehazing is crucial for remote sensing applications, we further evaluate our models on SateHaze1k [huang2020single] in Table VI. Our methods consistently achieve the best results across all subsets. In particular, PoolNet + UDP attains 26.20 dB, 28.26 dB, and 23.13 dB PSNR on the three subsets, outperforming both general dehazing models (e.g., FocalNet [cui2023focal], ConvIR-S [cui2024omni]) and remote sensing–oriented designs (e.g., Trinity-Net [chi2023trinity]). Notably, on the challenging Thick subset, our method achieves a gain of 1.44 dB over FocalNet, highlighting its strong generalization for remote sensing scenarios.
@@
To further demonstrate the superiority of our approach, we conduct experiments on the all-in-one settings [jiang2024survey]. The settings evaluate the ability of a single model to function as a universal and versatile restorer. In our study, we adopt the more challenging five-task benchmark, which jointly considers Gaussian denoising, deraining, dehazing, motion deblurring, and low-light enhancement, thus requiring stronger generalization from a single set of model weights. As shown in Table VII, both PromptIR + UDP and AdaIR + UDP achieve substantial improvements in dehazing and low-light enhancement, with PromptIR + UDP yields further gains in deraining and AdaIR + UDP further boosting performance in motion deblurring. This can be largely attributed to the integration of depth-based priors, which promote a more nuanced understanding of global atmospheric illumination and fine-grained object structures. Although originally introduced to enhance dehazing, such priors also strengthen the model’s ability to recover structural fidelity in motion deblurring and illumination details in low-light enhancement. Consequently, AdaIR + UDP surpasses its original counterpart and establishes a new benchmark for all-in-one image restoration.
@@
TABLE VII: Performance comparison of different methods on five challenging benchmarks. Denoising results are reported at $\sigma=25$. The baseline results are in blue. Best results are highlighted.
@@
  Method Dehazing on SOTS[li2018benchmarking] Deraining on Rain100L[yang2019joint] Denoising on BSD68[martin2001database] Deblurring on GoPro[nah2017deep] Low-Light on LOL[wei2018deep] Average
@@
  (TIP’23) DehazeFormer [song2023vision] 25.31 0.937 33.68 0.954 30.89 0.880 25.93 0.785 21.31 0.819 27.42 0.875
(ICCV’23) Retinexformer [cai2023retinexformer] 24.81 0.933 32.68 0.940 30.84 0.880 25.09 0.779 22.76 0.863 27.24 0.873
(ICCVW’21) SwinIR [liang2021swinir] 21.50 0.891 30.78 0.923 30.59 0.868 24.52 0.773 17.81 0.723 25.04 0.835
(CVPR’22) Restormer [zamir2022restormer] 24.09 0.927 34.81 0.960 31.49 0.884 27.22 0.829 20.41 0.806 27.60 0.881
(TPAMI’23) FSNet [cui2023image] 25.53 0.943 36.07 0.968 31.33 0.883 28.32 0.869 22.29 0.829 28.71 0.898
(CVPR’22) AirNet [li2022all] 21.04 0.884 32.98 0.951 30.91 0.882 24.35 0.781 18.18 0.735 25.49 0.846
(NeurIPS’23) PromptIR [potlapalli2023promptir] 26.54 0.949 36.37 0.970 31.47 0.886 28.71 0.881 22.68 0.832 29.15 0.904
(ICLR’25) AdaIR [cui2025adair] 30.53 0.978 38.02 0.981 31.35 0.889 28.12 0.858 23.00 0.845 30.20 0.910
(ICLR’25) $\text{DCPT}_{\text{PromptIR}}$ [hu2025universal] 30.72 0.977 37.32 0.978 31.32 0.885 28.84 0.877 23.35 0.840 30.31 0.911
(TPAMI’25) DA-RCOT [tang2025degradation] 30.96 0.975 37.87 0.980 31.23 0.888 28.68 0.872 23.25 0.836 30.40 0.911
(TIP’25) Perceive-IR [zhang2025perceive] 28.19 0.964 37.25 0.977 31.44 0.887 29.46 0.886 22.88 0.833 29.84 0.909
(TIP’25) Pool-AIO [cui2025exploring] 30.25 0.977 37.85 0.981 31.24 0.887 27.66 0.844 22.66 0.841 29.93 0.906
(TIP’25) $\text{DPPD}_{\text{PromptIR}}$ [wu2025learning] 30.31 0.980 37.32 0.980 31.33 0.885 28.74 0.875 22.73 0.846 30.09 0.913
(CVPR’25) VLU-Net [zeng2025vision] 30.84 0.980 38.54 0.982 31.43 0.891 27.46 0.840 22.29 0.833 30.11 0.905
@@
TABLE VIII: Ablation analysis on different variants of our method, mainly including two aspects: fusion stages and attention types. Here, all ablation models adopt the backbone of FSNet [cui2023image].
@@
V-D Ablation Study
We conduct ablation studies to demonstrate the effectiveness of the proposed UDPNet by training and testing the models on the Haze4K dataset [liu2021synthetic]. The model is trained for 1000 epochs using FSNet [cui2023image] as the baseline.
V-D1 Contribution of the Depth Map
The first step in our analysis is to justify the integration of depth information. Table IX shows that the baseline FSNet achieves 34.12 dB PSNR. By incorporating depth maps extracted using the DepthAnything-V2-small variant together with our proposed UDP, the model’s performance is improved to 34.69 dB, while only introducing 0.32 M parameters. Although the gain at this stage is relatively modest, it demonstrates the effectiveness of depth priors. Using the base and large variants of DepthAnything V2 further increases performance, reaching 34.87 dB and a peak of 35.31 dB, respectively. In contrast, when the depth map is replaced with a 128-level gray map, the performance drops to 34.08 dB, highlighting that the performance improvement comes from accurate depth information rather than the increase in parameter count. Overall, these results confirm that accurate depth estimation is a key factor in image dehazing.
TABLE IX: Comparison of different depth maps. All depth variants have an identical parameter increase of only +0.32M.
@@ V-D1 Contribution of the Depth Map
Figure 12: Performance and efficiency variations with our UDP on the Haze4K dataset, compared with baseline methods [cui2023image, cui2024revitalizing, cui2025exploring].
V-D2 Effectiveness of each module
We perform the breakdown ablations by applying our modules to the baseline successively. The number of heads in the multi-head attention is set to 1, while other configurations remain identical to the final dehazing model. As shown in Table VIII, directly concatenating refined depth maps and RGB images at the network head (row (b)) significantly improves performance compared to the baseline (row (a)), reaching 34.78 dB PSNR. This demonstrates the substantial benefit of introducing depth maps. By contrast, replacing simple concatenation with the proposed DGAM (row (c)) yields further gains with 34.96 dB PSNR, verifying that adaptively weighting depth cues through channel attention is more effective than naive feature stacking.
We further explore two variants of DPFM: channel cross-attention (CCA) and spatial cross-attention (SCA). Results in rows (d–g) show that both variants outperform DGAM, indicating the benefit of explicitly modeling cross-modal dependencies. Notably, using RGB features as the query (CCArgb, SCArgb) consistently outperforms the depth-as-query counterparts, suggesting that RGB semantics provide a more reliable guidance signal. Moreover, spatial cross-attention (SCA) shows slightly higher gains than channel cross-attention (CCA), suggesting that spatial correspondences play a crucial role in depth–RGB fusion. Across all cases, the best performance 35.22 dB PSNR is achieved in row (l), where both SCAdepth and SCArgb are combined in the encoder, highlighting the complementarity of depth and RGB at the spatial level and confirming the necessity of deliberate design.
@@ V-D2 Effectiveness of each module
V-D3 Position of DPFM
We also investigate inserting DPFM into different network stages. Placing DPFM in the encoder (rows (d–g)) yields more notable gains in Table VIII, since incorporating depth priors at early stages enhances semantic abstraction and provides stronger guidance for subsequent layers. By contrast, applying DPFM to the decoder (rows (h–k)) still brings consistent improvements, but the benefits are relatively smaller. This can be attributed to the fact that depth maps mainly contain low-frequency structural cues, which may interfere with the reconstruction of high-frequency image details when introduced in the decoder.
V-D4 Impact of the number of heads in DPFM
@@ V-D5 Generalization of our proposed UDP
To evaluate the adaptability of the proposed UDP, we integrate it into different baseline models, including FSNet [cui2023image], ConvIR [cui2024revitalizing], PoolNet [cui2025exploring], PromptIR [potlapalli2023promptir] and AdaIR [cui2025adair]. Each modified network is trained and tested under the corresponding experimental settings. The results demonstrate that UDP can be applied to a variety of multi-scale image restoration networks, while effectively enhance the performance of all-in-one image restoration models, yielding improvements in tasks such as dehazing, deblurring, deraining, and low-light enhancement.
V-E Limitations
@@
Our work introduces an effective method for image dehazing by leveraging depth-based priors, which significantly enhance adaptability and performance through plug-and-play modules. By integrating depth features, our approach lays the foundation for developing more robust and versatile dehazing models. However, the UDP has several limitations that open avenues for future research. One limitation is the reliance on depth priors extracted from degraded images, which may introduce inaccuracies due to noise or occlusions in the depth estimation process. These inaccuracies could lead to structural distortions or loss of fine details in the restored image. Furthermore, the dependence on DepthAnything V2 priors during inference introduces computational inefficiency, as the external model requires additional resources for depth estimation, increasing inference time. Future work could explore alternative strategies for real-time depth inference or distilling depth priors from depth estimation models into image restoration models. While the UDP framework demonstrates promising results across various restoration tasks, the potential of leveraging depth-based priors for general image restoration remains an exciting direction. Future studies will focus on refining adaptive strategies to improve performance in a broader range of scenarios, with particular attention to enhancing real-time performance.
VI Conclusion